A parallel plate capacitor is composed of two rectangular plates with length 5mm and width 3 mm. The thickness of the insulating material is 0.5 mm. Find the permittivity of the insulating material if the capacitance is 2 μF


1
Expert's answer
2021-05-20T18:24:30-0400

Length of rectangular plate, "L=5\space mm=5\times10^{-3}\space m"

Breadth of plate, "B=3\space mm=3\times10^{-3}\space m"

Area of plate, "A=15\times10^{-6}\space m^2"

Thickness of plate, "d=0.5\space mm=0.5\times10^{-3}\space m"

Capacitance, "C=2\space\mu F=2\times10^{-6}\space F"

Let permittivity of plate be "\epsilon"

Capacitance of parallel plate capacitor is given by

"C=\dfrac{\epsilon A}{d}"

Since, parallel plate capacitor is composed of two rectangular plates

"C=\dfrac{\epsilon A}{2d}"

"\epsilon=\dfrac{C\times2d}{A}"

"\epsilon=0.133\times10^{-3}\space C^2/N\space m^2"